# Black holes in the expanding Universe

### v2

@@ Abstract
The McVittie metric does not describe a physical black hole in an expanding universe because the curvature scalar and pressure at its event horizon are infinite. We show that extending this metric to an inhomogeneous scale factor, which depends on both the time and radial coordinate, removes those infinities by imposing the constancy of the Hubble parameter at the horizon. We also show that the Hubble parameters at the event horizons of all centrally symmetric black holes without accreting matter are equal to the same constant $H_{\textrm{hor}}=(\Lambda/3)^{1/2}$. Because of this equality and the equivalence to the Kottler metric near the horizon, black holes do not grow with the universe expansion. Black holes with accretion are described by this extension with different values of $H_{\textrm{hor}}$.
Introduction.
@@
McVittie metric.
The McVittie metric for a black hole, embedded in the flat, expanding FLRW universe, is given in the comoving spherical coordinates (with $c=1$) by McVittie
$ds^{2}=\Bigl{(}\frac{1-M}{1+M}\Bigr{)}^{2}d\tau^{2}-(1+M)^{4}a^{2}(dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta\,d\phi^{2}),\quad M(\tau,r)=\frac{Gm}{2ar},$ (1)
@@
where $H=\dot{a}/a$ is the Hubble parameter, the dot denotes differentiation with respect to $\tau$, and $\kappa=8\pi G$. The Ricci scalar is equal to $-12H^{2}-6\dot{H}(1+M)/(1-M)$.
@@
The event horizon of the black hole is given by $M=1$: $r=Gm/(2a)$. The McVittie metric (1) has a curvature singularity at the horizon because the Ricci scalar diverges there, unless $\dot{H}=0$. The pressure at the horizon also diverges, unless $\dot{H}=0$. However, that condition gives a constant Hubble parameter, describing an exponential expansion of the universe: $a\sim e^{H\tau}$, which is satisfied only for an empty universe with a positive cosmological constant. Therefore, the McVittie metric does not describe a physical black hole embedded in the FLRW spacetime because the curvature and pressure must be finite.
To remove the curvature and pressure singularities at the event horizon, one must impose there $\dot{H}=0$. However, astronomical observations show that the Hubble parameter has been changing in time, so $\dot{H}\neq 0$. Consequently, the scale factor at the event horizon must be different from the scale factor of the Universe as a whole. The scale factor must depend not only on the time $\tau$, but also on the radial coordinate $r$. Near a black hole, the scale factor is inhomogeneous. Far from a black hole, where $M\approx 0$, the scale factor and the Hubble parameter should asymptotically tend to functions of the time only, in accordance with the observed homogeneity and isotropy of the Universe on the large scale (the Cosmological Principle).
Tolman metric.
We consider the generalized metric (1) with the scale factor depending on both the time and radial coordinate. To determine the resulting field equations, we use the Tolman metric, which describes the geometry of a centrally symmetric gravitational field in spacetime filled with an ideal fluid LL2 ; Tol :
@@
where $\nu$, $\lambda$, and $\mu$ are functions of a time coordinate $\tau$ and a radial coordinate $R$. Applying coordinate transformations $\tau\rightarrow\tilde{\tau}(\tau)$ and $R\rightarrow\tilde{R}(R)$ does not change the form of the metric (4). The components of the Einstein tensor corresponding to (4) that do not vanish identically are LL2 ; Tol :
@@
In the isotropic spherical coordinates, the functions $\lambda$ and $\mu$ are related to one another: $e^{\mu}=R^{2}e^{\lambda}$. The Tolman metric (4) becomes
@@
[…]
@@
Field equations for generalized McVittie metric.
The metric (1) with an inhomogeneous scale factor $a(\tau,r)$ is the Tolman metric for $R=r$ and
$e^{\nu}=\Bigl{(}\frac{1-M}{1+M}\Bigr{)}^{2},\quad e^{\lambda}=(1+M)^{4}a^{2}.$ (11)
The corresponding Hubble parameter $H(\tau,r)$ is also inhomogeneous. We define
$F(\tau,r)=\frac{a^{\prime}}{a}.$ (12)
Accordingly, the following relations are satisfied:
$M^{\prime}=-M\Bigl{(}\frac{1}{r}+F\Bigr{)},\quad\dot{M}=-MH,\quad\nu^{\prime}=\frac{4M}{1-M^{2}}\Bigl{(}\frac{1}{r}+F\Bigr{)},\quad\lambda^{\prime}=-\frac{4M}{r(1+M)}+2F\frac{1-M}{1+M},$
$\nu^{\prime\prime}=-\frac{4M(1+M^{2})}{(1-M^{2})^{2}}\Bigl{(}\frac{1}{r}+F\Bigr{)}^{2}+\frac{4M}{1-M^{2}}\Bigl{(}-\frac{1}{r^{2}}+F^{\prime}\Bigr{)},\quad\dot{\nu}=\frac{4MH}{1-M^{2}},\quad\dot{\lambda}=2H\frac{1-M}{1+M},$
$\lambda^{\prime\prime}=\frac{4M}{(1+M)^{2}}\Bigl{(}\frac{1}{r}+F\Bigr{)}^{2}+\frac{4M}{r^{2}(1+M)}+2F^{\prime}\frac{1-M}{1+M},\quad\ddot{\lambda}=2\dot{H}\frac{1-M}{1+M}+\frac{4MH^{2}}{(1+M)^{2}},$
$\dot{\lambda}^{\prime}=2H^{\prime}\frac{1-M}{1+M}+\frac{4MH}{(1+M)^{2}}\Bigl{(}\frac{1}{r}+F\Bigr{)}.$ (13)
@@
$\kappa\epsilon=3H^{2}-\frac{1}{a^{2}(1+M)^{5}}\Bigl{(}\frac{4F}{r}+F^{2}(1+M)+2F^{\prime}(1-M)\Bigr{)},$
$\kappa p_{\textrm{r}}=-3H^{2}-2\dot{H}\frac{1+M}{1-M}+\frac{1}{a^{2}(1+M)^{4}}\Bigl{(}\frac{2F}{r}+F^{2}\Bigr{)},$
$\kappa p_{\textrm{t}}=-3H^{2}-2\dot{H}\frac{1+M}{1-M}+\frac{1}{a^{2}(1+M)^{5}}\Bigl{[}\frac{1+M^{2}}{1-M}\Bigl{(}\frac{F}{r}+F^{\prime}\Bigr{)}-\frac{2MF^{2}}{1+M}\Bigr{]},$
@@
The Ricci scalar is equal to $\kappa(p_{\textrm{r}}+2p_{\textrm{t}}-\epsilon)$.
These equations give the space and time dependence of the energy density, radial pressure, tangential pressure, and energy flux density near a black hole. They generalize the relations (2) to an inhomogeneous scale factor $a(\tau,r)$. The radial and tangential pressures are different; the pressure is anisotropic. The energy and both pressures are inhomogeneous; they depend on the radial coordinate $r$. Far from the black hole, the limits $r\to\infty$, $M\to 0$, $a(\tau,r)\to a(\tau)$, $a^{\prime}\to 0$, and $F\to 0$ reduce the field equations (14) to the Friedmann equations (3) for a homogeneous and isotropic fluid with $p_{\textrm{t}}=p_{\textrm{r}}$ and without energy flux $S=0$. Far from the black hole, the Universe is homogeneous and isotropic; as $r\to\infty$, the scale factor and the Hubble parameter tend to their values for the Universe on the large scale.
Constancy of Hubble parameter at an event horizon.
The first and fourth equation in (14) show that the energy density and the energy flux density are finite. Also, the fourth equation in (14) gives $S_{\textrm{hor}}=0$, so the energy flux corresponding to accretion of matter into a black hole vanishes at its event horizon, in accordance with gravitational time dilation. To ensure the regularity of the pressure and curvature for the metric (1) with an inhomogeneous scale factor $a(\tau,r)$ at the event horizon of a black hole, where $M=1$, the second and third equation in (14) at the horizon give
@@
Furthermore, the third equation in (14) at the horizon gives
$\Bigl{(}\frac{F}{r}+F^{\prime}\Bigr{)}_{\textrm{hor}}=0.$ (16)
The simplest solution satisfying both these conditions is the scale factor, for which the Hubble parameter near the horizon is constant, equal to $H_{\textrm{hor}}$ at the horizon, so $a\sim e^{H_{\textrm{hor}}\tau}$ near the horizon and thus $F_{\textrm{hor}}=H^{\prime}_{\textrm{hor}}=0$. Consequently, the first equation in (14) gives $\kappa\epsilon_{\textrm{hor}}=3H^{2}_{\textrm{hor}}$. If this solution represents a black hole that is not accreting any matter, then a constant $H_{\textrm{hor}}$ should be universal and related to the cosmological constant $\Lambda$ (which is the simplest explanation of the observed accelerating expansion of the Universe dark ):
$H_{\textrm{hor}}=(\Lambda/3)^{1/2},$ (17)
in accordance with $\kappa\epsilon=\Lambda$ in the Friedmann equations (3) in the absence of matter. Consequently, the Hubble parameters at the event horizons of all centrally symmetric black holes in the absence of accreting matter are equal. This relation also shows that the cosmological constant cannot be negative.

### v3

# Classical and Quantum Gravity 42 (6), 065017 (2025)

@@ Abstract
The McVittie metric does not describe a physical black hole in an expanding Universe because the curvature scalar and pressure at its event horizon are infinite. We show that extending this metric to an inhomogeneous scale factor, which depends on both the time and radial coordinate, removes those infinities by imposing at the horizon the constancy of the Hubble parameter and a particular constraint on the gradient of the scale factor. We consider a special case of this metric, and show that the Hubble parameters at the event horizons of all centrally symmetric black holes are equal to the same constant $H_{\textrm{hor}}=(\Lambda/3)^{1/2}$. Because of this equality and the equivalence to the Kottler metric near the horizon, black holes do not grow with the Universe expansion.
Keywords: black hole, expanding Universe, McVittie metric, event horizon, inhomogeneous scale factor, Hubble parameter, cosmological constant.
1. Introduction
@@
2. McVittie metric
The McVittie metric for a black hole, embedded in the flat, expanding FLRW Universe, is given in the comoving, isotropic spherical coordinates $\rho,\theta,\phi$ (with $c=1$) by McVittie
$ds^{2}=\Bigl(\frac{1-M}{1+M}\Bigr)^{2}d\tau^{2}-(1+M)^{4}a^{2}(d\rho^{2}+\rho^{2}d\theta^{2}+\rho^{2}\sin^{2}\theta\,d\phi^{2}),\quad M(\tau,\rho)=\frac{Gm}{2a\rho},$ (1)
@@ 2. McVittie metric
where $H(\tau)=\dot{a}/a=(da/d\tau)/a=d(\ln a)/d\tau$ is the Hubble parameter, the dot denotes differentiation with respect to $\tau$, and $\kappa=8\pi G$. The Ricci scalar is equal to $12H^{2}-6\dot{H}(1+M)/(1-M)$.
@@
3. Equivalency to Kottler metric
A coordinate transformation into a new radial coordinate $r$ Robertson :
$r=a(\tau)\rho(1+M)^{2},$ (4)
transforms the metric (1) into
$ds^{2}=\Bigl(1-\frac{2Gm}{r}-H^{2}r^{2}\Bigr)d\tau^{2}+\frac{2Hr}{(1-2Gm/r)^{1/2}}d\tau\,dr-\frac{dr^{2}}{1-2Gm/r}-r^{2}d\theta^{2}-r^{2}\sin^{2}\theta\,d\phi^{2}$
$=f\Bigl(d\tau+\frac{Hr}{(1-2Gm/r)^{1/2}f}dr\Bigr)^{2}-\frac{1}{f}dr^{2}-r^{2}d\theta^{2}-r^{2}\sin^{2}\theta\,d\phi^{2},$ (5)
where $f(\tau,r)=1-2Gm/r-H^{2}r^{2}$. All points having the same radial coordinate $r$ compose a sphere whose surface area is $4\pi r^{2}$.
If the Hubble parameter $H$ is constant, then a coordinate transformation into a new time coordinate $t$:
$t=\tau+H\int\frac{r}{(1-2Gm/r)^{1/2}f}dr,$ (6)
transforms the nondiagonal metric (5) into the diagonal Kottler (Schwarzschild–de Sitter) metric Kottler :
$ds^{2}=\Bigl(1-\frac{2Gm}{r}-\frac{1}{3}\Lambda r^{2}\Bigr)dt^{2}-\frac{dr^{2}}{1-2Gm/r-\Lambda r^{2}/3}-r^{2}d\theta^{2}-r^{2}\sin^{2}\theta\,d\phi^{2},$ (7)
where $\Lambda=3H^{2}$ is a positive cosmological constant. The cosmological constant is the simplest explanation of the observed accelerating expansion of the Universe dark . The Kottler metric describes a centrally symmetric black hole in a spacetime with a cosmological constant, reducing to the de Sitter metric deSitter in the absence of the black hole and to the Schwarzschild metric in the absence of the cosmological constant.
Because of the equivalency to the Kottler metric, the McVittie metric with a constant $H$ does not provide any new information. This case describes an exponential expansion of the Universe: $a\sim e^{H\tau}$, which is satisfied only for an empty Universe with a positive cosmological constant. Therefore, it does not describe a black hole in the real Universe with the presence of matter. Only the case with a time-varying Hubble parameter is of interest.
4. Event horizon
The event horizon of a black hole represented by the McVittie metric (1) is given by $M=1$:
$\rho_{\textrm{hor}}=\frac{Gm}{2a}.$ (8)
Because of the term $\dot{H}/(1-M)$ in (2), the pressure at the horizon diverges at the horizon, unless $\dot{H}=0$ (constant $H$). This metric has a curvature singularity at the horizon because the Ricci scalar also diverges there, unless $\dot{H}=0$. The case with a constant $H$ is described by the Kottler metric (7), for which the radius $r_{\textrm{hor}}$ of the event horizon of a black hole is constant and given by a cubic equation $1-2Gm/r_{\textrm{hor}}-\Lambda r^{2}_{\textrm{hor}}/3=0$. However, it does not describe a black hole in the real Universe, which has a time-varying Hubble parameter. Therefore, the McVittie metric does not describe a physical black hole embedded in the FLRW spacetime because the curvature and pressure must be finite Nolan ; singular .
To remove the curvature and pressure singularities at the event horizon, one must impose there $\dot{H}=0$. However, astronomical observations show that the Hubble parameter is changing in time, so $\dot{H}\neq 0$. Consequently, the scale factor at the event horizon must be different from the scale factor of the Universe as a whole. The scale factor must depend not only on the time $\tau$, but also on the radial coordinate $\rho$. Near a black hole, the scale factor is inhomogeneous. Far from a black hole, where $M\approx 0$, the scale factor and the Hubble parameter should asymptotically tend to functions of the time only, in accordance with the observed homogeneity and isotropy of the Universe on the large scale (the Cosmological Principle).
5. Tolman metric
We consider the generalized metric (1) with the scale factor depending on both the time and radial coordinate. To determine the resulting field equations, we use the Tolman metric, which describes the geometry of a centrally symmetric gravitational field in spacetime filled with an ideal fluid Tolman ; LL2 :
@@ 5. Tolman metric
where $\nu$, $\lambda$, and $\mu$ are functions of a time coordinate $\tau$ and a radial coordinate $R$. Applying coordinate transformations $\tau\rightarrow\tilde{\tau}(\tau)$ and $R\rightarrow\tilde{R}(R)$ does not change the form of the metric (9). The components of the Einstein tensor corresponding to (9) that do not vanish identically are Tolman ; LL2 :
@@
In the isotropic spherical coordinates, the functions $\lambda$ and $\mu$ are related to each other by a condition $e^{\mu}=R^{2}e^{\lambda}$. The Tolman metric (9) becomes
@@
Accordingly, $\mu=2\ln R+\lambda$, $\mu^{\prime}=2/R+\lambda^{\prime}$, $\mu^{\prime\prime}=-2/R^{2}+\lambda^{\prime\prime}$, $\dot{\mu}=\dot{\lambda}$, and $\ddot{\mu}=\ddot{\lambda}$. The Einstein equations (11) for the components of the Einstein tensor (10) reduce to
@@
6. Field equations for generalized McVittie metric
The metric (1) with an inhomogeneous scale factor $a(\tau,\rho)$:
$ds^{2}=\Bigl(\frac{1-M}{1+M}\Bigr)^{2}d\tau^{2}-(1+M)^{4}a^{2}(\tau,\rho)(d\rho^{2}+\rho^{2}d\theta^{2}+\rho^{2}\sin^{2}\theta\,d\phi^{2}),\quad M(\tau,\rho)=\frac{Gm}{2a(\tau,\rho)\rho},$ (16)
is a special case of the Tolman metric (13) for
$R=\rho,\quad e^{\nu}=\Bigl(\frac{1-M}{1+M}\Bigr)^{2},\quad e^{\lambda}=(1+M)^{4}a^{2}.$ (17)
Because the scale factor depends on both $\tau$ and $\rho$, to define the Hubble parameter as a partial derivative $\partial(\ln a)/\partial\tau$, it must be specified which radial coordinate is regarded as constant; the partial derivatives $\partial(\ln a)/\partial\tau]_{\rho}$ and $\partial(\ln a)/\partial\tau]_{r}$ with $r=a\rho(1+M)^{2}$ (4) are different. We define the Hubble parameter (and the dot in general) as the partial derivative with respect to time at a constant isotropic radial coordinate $\rho$:
$H(\tau,\rho)=\Bigl(\frac{\dot{a}}{a}\Bigr)_{\rho}=\frac{1}{a}\Bigl(\frac{\partial a}{\partial\tau}\Bigr)_{\rho}=\Bigl(\frac{\partial(\ln a)}{\partial\tau}\Bigr)_{\rho}.$ (18)
It is inhomogeneous like the scale factor. We also define
$F(\tau,\rho)=\Bigl(\frac{a^{\prime}}{a}\Bigr)_{\tau}=\frac{1}{a}\Bigl(\frac{\partial a}{\partial\rho}\Bigr)_{\tau}=\Bigl(\frac{\partial(\ln a)}{\partial\rho}\Bigr)_{\tau},$ (19)
which characterizes the gradient of the scale factor. Similarly, $\dot{H}=(\partial H/\partial\tau)_{\rho}$ and $H^{\prime}=(\partial H/\partial\rho)_{\tau}$. The prime denotes partial differentiation with respect to the isotropic radial coordinate $\rho$ at a constant time.
The following relations are satisfied:
$M^{\prime}=-M\Bigl(\frac{1}{\rho}+F\Bigr),\quad\dot{M}=-MH,\quad\nu^{\prime}=\frac{4M}{1-M^{2}}\Bigl(\frac{1}{\rho}+F\Bigr),\quad\lambda^{\prime}=-\frac{4M}{\rho(1+M)}+2F\frac{1-M}{1+M},$
$\nu^{\prime\prime}=-\frac{4M(1+M^{2})}{(1-M^{2})^{2}}\Bigl(\frac{1}{\rho}+F\Bigr)^{2}+\frac{4M}{1-M^{2}}\Bigl(-\frac{1}{\rho^{2}}+F^{\prime}\Bigr),\quad\dot{\nu}=\frac{4MH}{1-M^{2}},\quad\dot{\lambda}=2H\frac{1-M}{1+M},$
$\lambda^{\prime\prime}=\frac{4M}{(1+M)^{2}}\Bigl(\frac{1}{\rho}+F\Bigr)^{2}+\frac{4M}{\rho^{2}(1+M)}+2F^{\prime}\frac{1-M}{1+M},\quad\ddot{\lambda}=2\dot{H}\frac{1-M}{1+M}+\frac{4MH^{2}}{(1+M)^{2}},$
$\dot{\lambda}^{\prime}=2H^{\prime}\frac{1-M}{1+M}+\frac{4MH}{(1+M)^{2}}\Bigl(\frac{1}{\rho}+F\Bigr).$ (20)
@@ 6. Field equations for generalized McVittie metric
$\kappa\epsilon=3H^{2}-\frac{1}{a^{2}(1+M)^{5}}\Bigl(\frac{4F}{\rho}+F^{2}(1+M)+2F^{\prime}(1-M)\Bigr),$ (21)
$\kappa p_{\textrm{r}}=-3H^{2}-2\dot{H}\frac{1+M}{1-M}+\frac{1}{a^{2}(1+M)^{4}}\Bigl(\frac{2F}{\rho}+F^{2}\Bigr),$ (22)
$\kappa p_{\textrm{t}}=-3H^{2}-2\dot{H}\frac{1+M}{1-M}+\frac{1}{a^{2}(1+M)^{5}}\Bigl[\frac{1+M^{2}}{1-M}\Bigl(\frac{F}{\rho}+F^{\prime}\Bigr)-\frac{2MF^{2}}{1+M}\Bigr],$ (23)
@@
The Ricci scalar is equal to $\kappa(p_{\textrm{r}}+2p_{\textrm{t}}-\epsilon)$. These equations give the space and time dependence of the energy density, radial pressure, tangential pressure, and energy flux density near a black hole, described by the inhomogeneous McVittie metric. They generalize the relations (2) to an inhomogeneous scale factor $a(\tau,\rho)$. The radial and tangential pressures are different; the pressure is anisotropic. The energy and both pressures are inhomogeneous; they depend on the radial coordinate $\rho$.
7. Regularity at an event horizon
A singular behavior of the metric (16) at the event horizon of a black hole, where $M=1$, is a result of a singular behavior of $\nu^{\prime}$, $\nu^{\prime\prime}$, $\dot{\nu}$, and $e^{-\nu}$. To ensure the regularity of the pressure and curvature for this metric at the horizon, equations (22) and (23) give
@@ 7. Regularity at an event horizon
Furthermore, to ensure the regularity at the horizon, equation (23) gives $(F/\rho+F^{\prime})_{\textrm{hor}}=0$, which constrains the gradient of the scale factor. Let us consider a special case, in which the above relation is satisfied for an arbitrary $\rho$:
$\frac{F}{\rho}+F^{\prime}=0,\quad F=\frac{A(\tau)}{\rho},$ (26)
where $A(\tau)$ is some function of the time coordinate. Consequently, $(\ln a)^{\prime}=A(\tau)/\rho$, which integrated with respect to $\rho$ yields $\ln a=A(\tau)\ln\rho+\ln B(\tau)$, where $B(\tau)$ is some function of the time coordinate. The scale factor is therefore given by
$a(\tau,\rho)=B(\tau)\rho^{A(\tau)}.$ (27)
Accordingly, $\dot{a}=\dot{B}\rho^{A}+B\dot{A}\rho^{A}\ln\rho$, which gives
$H=\frac{\dot{B}}{B}+\dot{A}\ln\rho=(\ln B)^{\cdot}+\dot{A}\ln\rho,\quad H^{\prime}=\frac{\dot{A}}{\rho},\quad\dot{H}=(\ln B)^{\cdot\cdot}+\ddot{A}\ln\rho,\quad\dot{H}^{\prime}=\frac{\ddot{A}}{\rho}.$ (28)
At the horizon of a black hole, the relations (29) and (27) give
$\frac{Gm}{2\rho}=B\rho^{A},\quad\rho=\Bigl(\frac{Gm}{2B}\Bigr)^{1/(1+A)}.$ (29)
The regularity condition (25) with $\dot{H}$ given by (28) gives therefore
$(\ln B)^{\cdot\cdot}+\frac{\ddot{A}}{1+A}\ln\Bigl(\frac{Gm}{2B}\Bigr)=0.$ (30)
This equation relates the functions $A(\tau)$ and $B(\tau)$ in the scale factor (27). The metric (16) with this scale factor is regular at the horizon if the limit $\lim_{M\to 1}\dot{H}/(1-M)$ of the term $\dot{H}/(1-M)$ in equations (22) and (23) is finite. This term is indeterminate, of the form $0/0$. Resolving the indeterminacy by L’Hôpital’s rule and using (20), (26), (28), and (30), we obtain
$\lim_{M\to 1}\frac{\dot{H}}{1-M}=-\lim_{M\to 1}\frac{\dot{H}^{\prime}}{M^{\prime}}=\lim_{M\to 1}\frac{\ddot{A}/\rho}{M(1/\rho+F)}=\lim_{M\to 1}\frac{\ddot{A}}{1+\rho F}=\lim_{M\to 1}\frac{\ddot{A}}{1+A}=-\frac{(\ln B)^{\cdot\cdot}}{\ln(Gm/2B)}.$ (31)
In the scale factor (27), the function $B(\tau)$ describes the background scale factor, which exists in the absence of a black hole. Therefore, $(\ln B)^{\cdot\cdot}$ is finite. Also, $B\gg Gm$ (the Universe is much larger than the Schwarzschild radius of a black hole). Consequently, the limit (31) is finite and the metric (16) is regular at the horizon.
The function $A(\tau)$ describes the variation of the scale factor (27) with the radial coordinate $\rho$. It must satisfy
$|A(\tau)|\ll 1,$ (32)
so that the scale factor is approximately homogeneous, and $\lim_{\tau\to\infty}A=0$, so the Universe tends to a homogeneous form: $a=B(\tau)$. The function (27) would be exact for all values of radial coordinate $\rho$ if the black hole considered were the only body in the Universe. Therefore, its application is limited to the values of $\rho$, which are not too large, so the effects of other bodies on the metric can be averaged as the background scale factor. In the absence of a black hole, $m=0$ gives $A=0$, in accordance with (30), and the scale factor (27) is equal to its background form $a=B(\tau)$.
The energy density, radial pressure, tangential pressure, and energy flux density are determined by the field equations (21), (22), (23), and (24), with $F$ given by (26) and $H$ given by (28). They depend on the two functions $A(\tau)$ and $B(\tau)$, related to each other by (30). They are finite at the horizon because the functions $F$ and $H$, their derivatives $F^{\prime}$ and $H^{\prime}$, and the limit (31) are finite at the horizon. The energy flux density in (24) at the horizon is zero, which is consistent with the infinite gravitational time dilation of accretion of matter, represented by the background energy density and pressure, onto a black hole described by the metric (16). Because of this dilation, the mass $m$ in the metric (16) remains constant. For a more complete analysis, one must investigate whether the energy density and pressure determined by the field equations (21), (22), (23), and (24) can represent the types of matter and fields existing in Nature.
Far from a black hole, $F\to 0$, $F^{\prime}\to 0$, and $H^{\prime}\to 0$, so the field equations (21), (22), (23), and (24) reduce to equations (2) with $M\to 0$ (equivalent to the Friedmann equations (3) for $A\to 0$) for a nearly homogeneous and isotropic fluid with $p_{\textrm{t}}=p_{\textrm{r}}$ and without energy flux, $S=0$. The condition (32) guarantees that the metric (16) with the condition (26) can describe a black hole embedded in the Universe, which on the large scale is nearly homogeneous and isotropic.
The above analysis describes the special case of the metric, represented by the condition (26) on the gradient of the scale factor. In a more general solution, which can include a black hole accreting matter, for which the mass is a function $m(\tau,\rho)$, that condition must be satisfied only at the horizon (to ensure the regularity). More generally, the scale factor has the form (27) only near a black hole. Far from the black hole, the scale factor should not vary with $\rho$, but instead tend to $B(\tau)$ as $\rho\to\infty$, in accordance with the homogeneity and isotropy of the Universe on the large scale.
8. Constancy of Hubble parameter at an event horizon
The condition (25) means that the Hubble parameter at the event horizon of a black hole is constant in time. Using (28), (29), and (30), it is equal to
$H_{\textrm{hor}}=\lim_{M\to 1}[(\ln B)^{\cdot}+\dot{A}\ln\rho]=(\ln B)^{\cdot}+\frac{\dot{A}}{1+A}\ln\Bigl(\frac{Gm}{2B}\Bigr)=(\ln B)^{\cdot}-\frac{\dot{A}}{\ddot{A}}(\ln B)^{\cdot\cdot}.$ (33)
This constant value is equal to its value at $\tau\to\infty$, given by the asymptotic values $A\to 0$ and $B\to e^{H_{0}\tau}$, where $H_{0}=(\Lambda/3)^{1/2}$ (for $\Lambda\geq 0$), giving $\ln B\to H_{0}\tau$, $(\ln B)^{\cdot}\to H_{0}$, and $(\ln B)^{\cdot\cdot}\to 0$. Consequently, the Hubble parameter at an event horizon is equal to
$H_{\textrm{hor}}=\Bigl(\frac{\Lambda}{3}\Bigr)^{1/2},$ (34)
independently of the mass of a black hole.
Because the value (34) is a limit (33), it is valid not only in the special case of the metric, represented by the condition (26), but also in a more general solution, in which that condition is satisfied only at the horizon (to ensure the regularity). Therefore, the Hubble parameters at the event horizons of all centrally symmetric black holes are equal.